OpenStudy (anonymous):

Lori has determined a function f(x) that shows the exponential growth of the number of jeans Tim owns each year. Explain how the f−1^(x) can be found and what f−1^(250) means. I need help , i want to be taught step by step. I even prepared my own function to use. The function I chose to make for this problem is f(x)=5^x

OpenStudy (anonymous):

@mathmate can you help walk me through this please?

OpenStudy (domebotnos):

the x is 250?? we know that f(x) is y so have u found why yet??

OpenStudy (anonymous):

this is how far I've gone f(x)=5^x y=5^x x=5^y log5X= log5 5^y

OpenStudy (anonymous):

@domebotnos

OpenStudy (anonymous):

@doc.brown

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

You were almost there! y=5^x exchange x and y x=5^y \(log_5 x = y\) by the definition of log or \(f^{-1}(x)=y=log_5x\) so \(f^{-1}(250)=y=log_5 250=log_5 5^3(2)=log_5 5^3+log_5 2=3+log_52\) using properties of logarihms.

OpenStudy (anonymous):

@mmathmate thanks a lot!~

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

@mathmate wait, can you explain the middles step a little clearer for me?

OpenStudy (mathmate):

\(f^{-1}(250)=y=log_5 250\).....from definition of inverse, substitute x=250 \(=log_5 5^3(2)\).................factorization of 250 \(=log_5 5^3+log_5 2\)..........log ab = log a + log b \(=3log_5 5 + log_5 2\) \(=3+log_52\) ...............from definition of logarithm \(log_a b=X \Leftrightarrow b=a^X\) so \( log_5 5 \Leftrightarrow 5=5^1\) or \(log_5 5=1\)

OpenStudy (anonymous):

@mathmate how did you get the 3 in the exponent?

OpenStudy (mathmate):

in the factorization, \(250 = 125*2 = 5*5*5 * 2 = 5^3 * 2\) if that's the part you were querying.

OpenStudy (anonymous):

@mathmate oh i get that part now, thanks. But how did you get to =3+log5 2?

OpenStudy (mathmate):

The definition of log is \(log_a b=X \Leftrightarrow b=a^X\) That means the two expressions on each side of the double arrow are equivalent (for all values of a,b,X within the domain of log). So if X=1, and a=2, then b=2^1=2, or \(log_2 2=1\) Similarly, if X=1, and a=5, then b=5^1=5, or \(log_5 5=1\) In fact, in general \(log_a a = 1\) for all values of a>0. That is how \(log_5 5\) was evaluated to 1, which gives \(3log_5 5=3\)

OpenStudy (anonymous):

@mathmate can you please simplify that for me? haha

OpenStudy (mathmate):

Let's go line by line. Are you familiar with the definition of log, shown on the second line of the last post?

OpenStudy (mathmate):

The definition of log means that if \(log_a b=X\) then it implies that \(b=a^X\) AND if \(b=a^X\) the it implies that \(log_a b=X\).

OpenStudy (mathmate):

* for all all valid values of a, b and X.

OpenStudy (anonymous):

@mathmate yes i understood this, whats after the equal sign is whats going to be exponent of the base correct? and the argument will be on the other side of the equal sign, right?

OpenStudy (mathmate):

So if 5^1 =5, what is \(log_5 5\) straight from the definition of log?

OpenStudy (anonymous):

@mathmate can you rephrase that question?

OpenStudy (mathmate):

Given 5^1=5 (which you already know), can you find what is X=\(log_5 5\) using the definition of log: \(log_a b=X \Leftrightarrow b=a^X\)

OpenStudy (mathmate):

hint: put a=b=5

OpenStudy (anonymous):

do you mean log5 5 =x to 5=5^x ?

OpenStudy (mathmate):

Yes, what do you find for x?

OpenStudy (anonymous):

is this where the factoring comes in and i choose the value of x to be 3?

OpenStudy (mathmate):

You cannot \(choose\) x if 5^x=5, you need to \(solve\) for x because there is only one correct value of x. Hints: 5^0=1 5^1=5 5^2=25 5^3=125

OpenStudy (anonymous):

is it 125?

OpenStudy (anonymous):

wait, its 2

OpenStudy (mathmate):

5^125\(\ne\)5 !!!

OpenStudy (mathmate):

\(5^2\ne5\) either.

OpenStudy (anonymous):

its 3 because 5*5*5=125 correct?

OpenStudy (mathmate):

The question we are trying to solve is \(log_5 5=?\) We are trying to solve an equivalent problem (using the definition of logarithm): 5^x=5 so 5*5*5=125 is not relevant. \(5^x=5*....=5\) is relevant.

OpenStudy (mathmate):

"Five to what power will give you five?" is the equivalent problem we are trying to solve.

OpenStudy (anonymous):

is it 1?

OpenStudy (mathmate):

Exactly, that's how we got one in the last step. Remember, Hints: 5^0=1 5^1=5 5^2=25 5^3=125 ...

OpenStudy (mathmate):

so just a check-up, what is \(log_5 25\)?

OpenStudy (anonymous):

17.47?

OpenStudy (mathmate):

solve the equivalent problem for x 5^x=25

OpenStudy (anonymous):

5^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

right*

OpenStudy (mathmate):

so what is x?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

?

OpenStudy (mathmate):

right! Let's put it together: when you are given a problem to evaluate \(\log_5 125=?\) 1. Assume \(\log_5 125=x\) and try to find x by the following. 2. formulate the equivalent problem using the definition of logarithm \(log_5 125=x\) is the same problem as \(5^x=125\) 3. Solve for x in the last equation, using trial and error, or using the calculator. (x=3) 4. Go back to the original problem and state that \(log_5 125=3\).

OpenStudy (mathmate):

Would that be a good summary to follow? @cupcakes123

OpenStudy (mathmate):

right! Let's put it together: when you are given a problem to evaluate \(\log_5 125=?\) 1. Assume \(\log_5 125=x\) and try to find x by the following. 2. formulate the equivalent problem using the definition of logarithm \(log_5 125=x\) is the same problem as \(5^x=125\) 3. Solve for x in the last equation, using trial and error, or using the calculator. (x=3) 4. Go back to the original problem and state that \(log_5 125=3\).

OpenStudy (anonymous):

so it would be log5^3+log5^2=250?

OpenStudy (mathmate):

I am not sure where the expression is coming from, can you explain?

OpenStudy (anonymous):

is it log5 5^3 +log5 2

OpenStudy (mathmate):

Which problem are you solving? Can you please explain in detail?

OpenStudy (anonymous):

so f^-1(x) would be equal to that?

OpenStudy (anonymous):

im trying to say, does f^-1(250) is equal to log5 of 5^3?

OpenStudy (anonymous):

+log5 of 2

OpenStudy (mathmate):

\(f^{-1}(250)\) has already been solved above, and I extract the last line: \(f^{-1}(250)=y=log_5 250=log_5 (5^3\times 2)=log_5 5^3+log_5 2=3+log_52\) I have changed the parentheses to a multiplication symbol, hope that's more clear.

OpenStudy (anonymous):

so, since they have the same base, which is 5, they cancel out and i multiply the arguments 5^3 times 2?

OpenStudy (mathmate):

\(log_5 (125)=3\) because \(5^3=125\).

OpenStudy (mathmate):

I don't understand what "they" stands for in "they" have the same base.

OpenStudy (anonymous):

sorry lol i meant it has

OpenStudy (mathmate):

In the last line, do you understand every step, which I broke down step by step above. Are you sure you understand every step, and why?

OpenStudy (anonymous):

yes i understood that last line

OpenStudy (mathmate):

250=5^3 \(\times\)2. Since 2 is not an integer power of x, it has to remain as \(log_5 2\). Since 125=5^3, we can simplify \(log_5 125=3\) So \(f^{-1}(250)=log_5 250=log_5 125 + log_5 2=3 + log_5 2\)

OpenStudy (mathmate):

Remember log (ab) = log(a) + log(b), whatever the base of log.

OpenStudy (anonymous):

Thank you very much for putting up with me. Im really sorry you have a lot of patience lol thank you so mcuh for helping really appreciated. thank you

OpenStudy (mathmate):

Are you sure you have understood?

OpenStudy (anonymous):

yes, i understand now :)

OpenStudy (mathmate):

Very well! Good night! P:)